The time T required to do a job varies inversely as the number of people P working. It takes 5 hours for 7 bricklayers to build a park wall. How long will it take for 12 bricklayers to complete the job?
Answer by mananth(16946)

5 hours 7 brick layers
12 bricklayer = how many hours.
..
It will take less time.
7/5 = x/12
inverse proportion
= 7/12 * 5
=35/12
=2.91 hours